# Net Charge Accretion in Magnetized Kerr Black Holes

### v2

@@ Appendix C The large $\varepsilon$ limit of $b_{0}^{-}$
Though not directly relevant for the main claim of the present paper, here we explain another application of the universal limit $\varepsilon\to$-$\infty$, namely the (semi-)analytical computation of the upper bound on the absorption cross section of an attracted particle, $b_{0}^{-}$. The procedure implemented at finite $\varepsilon$ for finding the critical points of the normalization function $n$ (21) is directly applicable to the universal equation, and it yields the analogous set of equations (28), (29), (30), allowing for a numerical solution for the limiting value. The finite-$\varepsilon$ values calculated for $b_{0}^{-}$ show excellent convergence to this limiting value, in analogy to what was shown for $b_{1}^{-}$ in Fig. 2.
However, we found that this value may be argued for in an even more elegant manner. The normalization in the universal limit attains a remarkably simple form:
@@
When considering a particle as it crosses the equatorial plane, (46) becomes even simpler, and reduces to a quadratic polynomial in $b^{2}$. For any given $r>2$, the roots of the polynomial then define the range of $b$ values that would allow such a charge to cross the equatorial plane at that radius. The larger root attains a minimum value that corresponds to the maximal impact parameter for which the orbit energetically allows the particle to cross the surface defined by $r=2$. Since all particles with $b$ greater than the largest root of the polynomial are never allowed to cross this barrier, they are of course incapable of falling into the BH since $r_{+}\leq 2$ for all $\alpha$. This value, dependent only on $\alpha$, proves to be equal (for all $\alpha$) to the limiting value calculated by solving the three equations mentioned above, and may be presented simply by the following form Okunthesis:
$\lim_{\varepsilon\rightarrow\infty}\mathopen{}\mathclose{{\left(b_{0}^{-}}}\right)^{2}=\min_{r}\frac{r\Delta\mathopen{}\mathclose{{\left(r}}\right)+2a\sqrt{\Delta\mathopen{}\mathclose{{\left(r}}\right)}}{r-2}~.$ (47)
@@ Appendix D An example of the classification process of critical points
As discussed in Sec. IV, the critical points can be divided into two categories: equatorial points and non-equatorial points. For a critical point to be considered as relevant, the sign of the determinant of the Hessian matrix of $n_{\mathrm{u}}$ evaluated at this point must be negative. Furthermore, for equatorial points, the additional condition of $n_{\mathrm{u},rr}<0$ must be met. After the relevant points have been identified, we seek out the points which are locally disconnecting, i.e. $n_{\mathrm{u},b}>0$. Finally, to discern between remaining candidates, we identify as corresponding to $b_{0}$ the critical point whose $b$ value is smaller than $b_{h}$.
In this domain there are only four critical points. One of them is non-equatorial, while the other three are on the equatorial plane. While their actual $r,\theta$ and $b$ values change as a function of $\varepsilon$, their ordering and the properties relevant to the classification process remain unchanged throughout the entire domain. We find that for these values of $\alpha$ and $E$, the critical point corresponding to $b_{0}$ is always the single non-equatorial one.

### v3

@@ Appendix C The large $\varepsilon$ limit of $b_{0}^{-}$
Although not directly relevant for the main claim of this paper, here we present another application of the universal limit $\varepsilon\to$-$\infty$ (see Subsec. III.3) – namely, the (semi-)analytical computation of the upper bound $\pi\mathopen{}\mathclose{{\left(b_{0}^{\mathrm{univ}}}}\right)^{2}$ on the large-$\mathopen{}\mathclose{{\left|\varepsilon}}\right|$ absorption cross section of an attracted particle, where $b_{0}^{\mathrm{univ}}\equiv\lim_{\varepsilon\to-\infty}b_{0}^{-}(\varepsilon)$. The CQB procedure implemented (in Sec. IV) at finite $\varepsilon$ for finding the critical points of the normalization function $n$ (Eq. (25)) is directly applicable to the normalization condition (Eq. (20)) associated with the universal $\varepsilon\to$-$\infty$ limit. It yields a set of equations analogous to Eqs. (28), (29), (30), allowing for a direct computation of the limiting value $b_{0}^{\mathrm{univ}}$. The finite-$\varepsilon$ values calculated for $b_{0}^{-}$ (using the methods described in Sec. IV) show excellent convergence to this limiting value – in analogy to what was shown for $b_{1}^{-}$ in Fig. 2.
However, we found that the value of $b_{0}^{\mathrm{univ}}$ may be obtained through a computationally simpler procedure, as we now sketch.
The normalization in the universal limit attains a remarkably simple form:
@@
When considering a particle as it crosses the equatorial plane, Eq. (46) becomes even simpler. For any given value of $r>2$, this polynomial in $b$ has two real positive roots which we denote $b_{\mathrm{root},i}(r)$ (where $i=1,2$). These roots then dictate the range of $b$ values that would allow the particle to arrive at the equatorial plane at that $r$ value. Upon further investigation one finds that the larger root, $b_{\mathrm{root},2}(r)$, attains a minimum value at some $r\equiv r^{\mathrm{min}}>2$ – which we denote $b_{\mathrm{root},2}^{\mathrm{min}}$. Moreover, we find Okunthesis that this $b$ value actually corresponds to the maximal impact parameter for which the orbit is energetically allowed to cross the surface defined by $r=2$ (at any $\theta$). Consequently, every trajectory with $b>b_{\mathrm{root},2}^{\mathrm{min}}$ is obviously energetically disallowed to fall into the BH (since $r_{+}\leq 2$ for all $\alpha$). As it turns out, $b_{\mathrm{root},2}^{\mathrm{min}}$ proves to be equal to $b_{0}^{\mathrm{univ}}$ [^12] This equality reflects the fact that in the context of the universal limit, the critical point associated with $b_{0}^{\mathrm{univ}}$ is equatorial.. This value, which depends only on $\alpha$, may be presented compactly by:
$\mathopen{}\mathclose{{\left(b_{0}^{\mathrm{univ}}}}\right)^{2}=\min_{r}\frac{r\Delta+2\alpha\sqrt{\Delta}}{r-2}~,$ (47)
where, recall, $\Delta\equiv r^{2}-2r+\alpha^{2}$. Equation (47) allows us to directly obtain $b_{0}^{\mathrm{univ}}$ in a simple (semi-)analytical manner.
@@ Appendix D An example of the classification process of critical points
As discussed in Sec. IV, the critical points can be divided into two categories: equatorial points and non-equatorial points. For a critical point to be considered as relevant, the sign of the determinant of the Hessian matrix of $n_{\mathrm{u}}$ evaluated at this point must be negative. Furthermore, for equatorial points, the additional condition $n_{\mathrm{u},rr}<0$ must be met. After the relevant points have been identified, we seek out the points which are locally disconnecting, i.e. $n_{\mathrm{u},b}>0$. Finally, to discern between remaining candidates, we identify as corresponding to $b_{0}$ the critical point whose $b$ value is smaller than $b_{h}$.
In this domain there are only four critical points. One of them is non-equatorial, while the other three are on the equatorial plane. While their actual $r,\theta$ and $b$ values change as a function of $\varepsilon$, their ordering and the properties relevant to the classification process remain unchanged throughout the entire domain. We find that for these values of $\alpha$ and $E$, the critical point corresponding to $b_{0}$ is always the single non-equatorial one. A summary of the classification process is presented in Table 1.